OpenStudy (anonymous):

What is the missing side length of an irregular hexagon with side lengths eleven point seven centimeters, nine point four centimeters, twenty point three centimeters, fifteen centimeters and twenty-one point one centimeters, if the perimeter is equal to the perimeter of a regular pentagon with side length twenty-one centimeters?

OpenStudy (anonymous):

Ill give medals!

OpenStudy (anonymous):

six point five centimeters nine point three centimeters twenty-seven point five centimeters forty-eight point five centimeters

OpenStudy (anonymous):

Well, then regular pentagon has 5 equal sides of 21 cm. So, 21*5=105 Taking the irregular sides of the hexagon: 11.7+9.4+20.3+15+21.1=77.5 105-77.5=27.5 Ask me if you need more help.

OpenStudy (anonymous):

Where's my medal?

OpenStudy (anonymous):

There

OpenStudy (anonymous):

Thanks. :)